OpenStudy (anonymous):

PLEASE HELP NEED TO COMPLETE FAN f(t)=-16t^2+94+12 which of the following is a reasonable domain of the graph of the function when the basketball falls from its highest height to the ground 3<=x<=6 2<=x<=5 1<=x<=4 0<=x<=3

OpenStudy (anonymous):

please help

OpenStudy (alexwee123):

oh my did i hear the f word? XD before we start did you make a mistake in the equation? is there supposed to be a t next to the 94?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

will give medal

OpenStudy (alexwee123):

ok first we need to find out where the maximum point is the general equation to quadratics equations like this is y=Ax^2+bx+c we'd use x=-b/2a to find the x coordinate of the maximum point you following me?

OpenStudy (anonymous):

yes

OpenStudy (alexwee123):

ok ill set it up for you then you can do the simplification \[X=\frac{ -94 }{ 2\times-16 }\] solve for x please

OpenStudy (anonymous):

how?

OpenStudy (alexwee123):

you divide -94 with (2 times -16) you might need a calculator

OpenStudy (anonymous):

-2.9375?

OpenStudy (anonymous):

help

OpenStudy (alexwee123):

yes good job if we round the answer we'd get 3 now which of the answer choices contain 3?

OpenStudy (anonymous):

the first one?

OpenStudy (anonymous):

the last one

OpenStudy (alexwee123):

good job ok now which one seems to be the answer? remember* at x=3 is when the function is at its maximum

OpenStudy (anonymous):

I think the last one

OpenStudy (anonymous):

am I correct I think its the last one

OpenStudy (anonymous):

please I will be your fan can you help me

OpenStudy (alexwee123):

no if it's the last one then it's stating that at x=0 is the maximum and that it ends at x=3

OpenStudy (anonymous):

so the first one?

OpenStudy (alexwee123):

yes

OpenStudy (anonymous):

thanks I will be your fan

OpenStudy (anonymous):

can you help me with another question please